OpenStudy (anonymous):

consider the quadratic function f(x)=-2x(squared)+12x-10 , find coordinates of the vertex:(x,y)?

OpenStudy (tkhunny):

Completing the Square might be the first line of defense.

OpenStudy (anonymous):

\[f(x)=-2x^2+12x-10\]first coordinate of the vertex is \(-\frac{b}{2a}\) in your example it is \[-\frac{12}{2\times (-2)}=3\]

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by \(3\)

OpenStudy (anonymous):

Thank you for your help. I can probably figure it out from here. Thanks again.